Ans. During one aycle, an engine extracts 2.00 X 10 J of energy from a hot reservoir transfers 1.5 X 103 J to a cold reservoir. 2. and a. Find the thermal efficiency of the engine Ans. b. How much work does the engine do in one cycle? Ans. c. How much power does the engine generate if it goes through four cycles in 2.5 see?
Answer #1

(a)

QH QT 2.00 109 1.5 10 0.25 QH 2.00-103

(b)

ll. = QH-OT-(2.00-1.5) . 103 J = 0.5 × 103 J

(c)

W 4 0.5 10 P=-= W 800 W 2.5